Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A 6-sided die has 3 black sides and 3 white sides. If the

Expert replies
by BTGmoderatorDC » Sat Apr 27, 2019 5:43 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

A 6-sided die has 3 black sides and 3 white sides. If the die is thrown 4 times, what is the probability that, on at least one of the throws, the die will land with a black side up?


A. 1/16

B. 3/16

C. 1/2

D. 9/16

E. 15/16

OA E

Source: Princeton Review
Join the discussion
Source: — Problem Solving |

by Jay@ManhattanReview » Mon Apr 29, 2019 9:41 pm
BTGmoderatorDC wrote:A 6-sided die has 3 black sides and 3 white sides. If the die is thrown 4 times, what is the probability that, on at least one of the throws, the die will land with a black side up?

A. 1/16

B. 3/16

C. 1/2

D. 9/16

E. 15/16

OA E

Source: Princeton Review
Since the 6-sided die has 3 black sides and 3 white sides, i.e., an equal number of black and white sides, the probability of getting a black side up = 1/2

Instead of calculating the probability that, on at least one of the throws, the die will land with a black side up, let's calculate the probability that, on none of the throws, the die land with a black side up. Once we get this, we will deduct the result from 1 to get the required answer.

Probability that, on none of the throws, the die land with a black side up = 1/2 * 1/2 *1/2 *1/2 = 1/16

Thus,

Probability that, on at least one of the throws, the die will land with a black side up = 1 – 1/16 = 15/16

The correct answer: E

Hope this helps!

-Jay
_________________
Manhattan Review GMAT Prep

Locations: GRE Manhattan | TOEFL Prep Courses New York | IELTS Prep Courses Boston | Seattle ACT Tutoring | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

by swerve » Thu May 02, 2019 9:41 am
\(P\) of each black side \(= \frac{3}{6} = \frac{1}{2}\)

and \(P\) of each white side \(= \frac{3}{6}=\frac{1}{2}\)

so die will land with a black side \(= 1- \text{white side } P\)

\(1- \left( \frac{1}{2}\right)^4 = \frac{15}{16}\)
Join the discussion

by Scott@TargetTestPrep » Thu May 02, 2019 4:20 pm
BTGmoderatorDC wrote:A 6-sided die has 3 black sides and 3 white sides. If the die is thrown 4 times, what is the probability that, on at least one of the throws, the die will land with a black side up?


A. 1/16

B. 3/16

C. 1/2

D. 9/16

E. 15/16

OA E

Source: Princeton Review

We see that P(black) = 3/6 = 1/2 and P(white) = 3/6 = 1/2. We can use the formula:

P(at least one black face comes up in 4 throws) = 1 - P(no black faces come up in 4 throws)

In other words,

P(at least one black face comes up in 4 throws) = 1 - P(all white faces come up in 4 throws)

Therefore,

P(at least one black face comes up in 4 throws) = 1 - (1/2)^4 = 1 - 1/16 = 15/16.

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

by deloitte247 » Sat May 04, 2019 9:49 pm
$$\Pr obability\ of\ a\ throw\ showing\ white\ at\ the\ top=\frac{3}{6}=\frac{1}{2}$$
$$\Pr obability\ of\ a\ throw\ showing\ black\ at\ the\ top=\frac{3}{6}=\frac{1}{2}$$
If the die is thrown 4 times, the probability that, on at least one of the throws, the die will land with a black = 1 - (Pr. of the die due to land with a white face up at all 4 times) $$=1-\left(\frac{1}{2}\cdot\frac{1}{2}\cdot\frac{1}{2}\cdot\frac{1}{2}\right)$$
$$=1-\frac{1}{16}$$
$$=\frac{\left(16-1\right)}{16}$$
$$=\frac{15}{16}\ \ \ \ \ \ \ \ \ \left(option\ E\right)$$
Join the discussion